OpenStudy (anonymous):

What is the slope of the line passing through the points (–1, 3) and (4, –7)? A. 2 B. 3/4 C -4/3 D. -2

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

This would get a better answer if you post it in the MATH SECTION :)

OpenStudy (triciaal):

slope is the change in y/ corresponding change in x for the points (x1, y1) and (x2, y2) slope,m = (y2 - y1)/(x2 - x1) for this problem one set of values (x1, y1) and the other (x2, y2) be consistent and substitute

OpenStudy (triciaal):

m =( -7 -3)/(4 - -1) = -10/5 = -2 so D